OpenStudy (anonymous):

How is completing the square beneficial? I know how to complete the square, but I don't why it's helpful.

OpenStudy (amistre64):

completing a square allows us to reduce a question from x^2 into x .... so that x can be solved for

OpenStudy (amistre64):

y = x^2 -4x +7 completes into y = (x-2)^2 + 3 the completed square is simpler to solve for x y-3 = (x-2)^2 +- sqrt(y-3) = x-2 etc ...

OpenStudy (anonymous):

Can you help me on completing the square??

OpenStudy (amistre64):

do you have your own post to play on?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@amistre64 Thank you so much. :)

OpenStudy (amistre64):

youre welcome, and good luck :)